OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form. –x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3

OpenStudy (anonymous):

slope/gradient = -1 and given the point (-1,4) we can use the find an equation formula to find the actual equation y-y1 = m (x-x1) y - 4 = -1 (x--1) y-4 = -1 (x+1) y-4 = -x -1 y = -x -1 + 4 y = -x + 3 so the third one because it can be rearranged to be y = -x +3